Solve for x and explain:
log[5] x=-2
Hint: [] stands for the number is written is subscript!

Answer by nerdybill(5404)

From "log rules" we have (using your notation):
log[a] x = N means that a^N = x
where ^ represents exponent
.
So, for:
log[5] x=-2
x = 5^(-2)
x = 1/5^2
x = 1/25
.
Review log rules here:
http://www.mathwords.com/l/logarithm_rules.htm